To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bars included 2 tagged bears. What is the conservationists estimate of the size of the bear population?
Use a porportion to solve

Answer by stanbon(75887)

To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bars included 2 tagged bears. What is the conservationists estimate of the size of the bear population?
Use a porportion to solve
----
(pop #)/(orig # tagged) = (# caught include tagged)/(# tagged of those caught)
---------------------
(pop #)/50 = 100/2
---
pop# = 50*50 = 2500
---------------------------
Cheers,
Stan H.